July 13-15

– Four stores provided ~21 hours of luminosity
– Tevatron heat exchanger work continued
– MI LCW leak found and fixed
– Pbar LCW leak still a mystery
– Store 8891 aborted due to Tevatron vacuum valves
– MTA took beam
– Search for Tevatron ground fault halted when fault disappeared
– KRF6 and LRF4 problems held off shot setup for approximately six hours

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts